OpenStudy (anonymous):

what is the radius of a circle with an area of 9 pi

OpenStudy (anonymous):

2 pi 9 3 none of above

OpenStudy (anonymous):

\[A=\pi r^2 \Rightarrow 9 \pi = \pi r^2\] solve for r

OpenStudy (anonymous):

i know that... i dont have r

OpenStudy (anonymous):

yes, you're supposed to solve for r... \(9\pi=\pi r^2\) divide both sides by \(\pi\) and then take the square root fo each side... what do you get?

OpenStudy (anonymous):

i got .1444 9*3.14=3.14*r*r 9*3.14=28.26 28.26/2=14.13 r=14.13

OpenStudy (anonymous):

\[9\pi = \pi r^2 \Rightarrow 9 = r^2 \Rightarrow r = 3\]

OpenStudy (anonymous):

so 9*3.14 should = 3.14*r squared. the that is nine which equals r squared and r is 3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@zaibali.qasmi @ganeshie8 @mathmale (Moderator)

OpenStudy (mathmale):

If you have a value for the radius, r, now, recall that the area of a circle is pi*(radius)^2. Substitute your value for r. If the result is the same as you started with (9pi square units), then you know your answer (for the radius) is correct.

OpenStudy (anonymous):

use this realtion \[r=\sqrt{\frac{ A }{ \pi }}\] put values and solve..

OpenStudy (mathmale):

@zaibali.quasmi: Would you mind explaining to math-geek where this formula came from, and why you obtained it?

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

what if its like 4pi

OpenStudy (anonymous):

well.. see formula for area of circle is A=pi * r^2 => A/pi= r^2 =>r=sqrt A/pi if you see the math geek coments carefully..he used this formula..

OpenStudy (anonymous):

true...true